# Yuxia Fu  Zhizhen Zhang11footnotemark: 1  Yuqi Zhang  Zijian Wang  Zi Huang  Yadan Luo

### v1

@@ 1 Introduction
Vision-Language-Action (VLA) models [OpenVLA, pi0, openvla-oft, pi0.5, rt1, rt2, gr00t, spatialvla, smolvla, octo, tinyvla, magma, vlm2vla] have recently enabled robot agents to perform complex manipulation tasks by fine-tuning large vision-language models (VLMs) with millions of robotic demonstrations. By reformulating action learning as a language generation or policy decoding problem, VLA models inherit broad visual grounding and semantic understanding from VLMs, and have shown notable performance in single-task or single-embodiment settings. However, real-world generalist agents must support multiple skills, embodiments, and environments, requiring the ability to consolidate many independently fine-tuned VLA models into a single unified policy.
@@
A natural approach is model merging, which has proven effective in large language and vision models [TA, ties, iso-cts, wudi, dare, emr-merging, knots, robust-merging, pem-composition]. Those techniques can integrate multiple specialized models without joint retraining or revisiting their original datasets. Yet, when these approaches are applied to VLA experts fine-tuned on distinct manipulation tasks, the merged model exhibits near-zero success rate. This failure suggests that VLA finetuning induces structural specialization incompatible across tasks, which is rarely observed in conventional VLM merging.
@@
To uncover the root causes, as shown in Figure 1, we perform a fine-grained decomposition of trainable parameters across representative VLA architectures and identify two major sources of non-mergeability: First, LoRA updates within the VLM backbone diverge sharply across tasks. Each manipulation task reshapes the pretrained representation space to satisfy its own perception-control alignment. Direct averaging  [TA, knots] or sign-resolved merging [ties] thus reactivates irrelevant or even contradictory parameters, corrupting shared vision-language subspaces and degrading task-invariant semantics. Second, the train-from-scratch action decoders accumulate strong task-specific dependencies across blocks through self-attention feedback. This coupling spreads localized task information globally, breaking modularity and preventing compositional merging even under identical architectures and initialization.
@@
Built upon these findings, we introduce MergeVLA, a merge–oriented VLA architecture that preserves mergeability by design. When executing a specific task, MergeVLA applies sparsely activated LoRA adapters, implemented via task masks, to selectively activate the merged parameters contributing to task-relevant responses while suppressing those that mislead other tasks. Moreover, MergeVLA reconfigures the action expert to remove self-attention propagation, relying solely on cross-attention pathways. This eliminates the sustained accumulation of task dependence across blocks, allowing most layers to be effectively merged using simple weight averaging. Due to these architectural modifications, it surprisingly shows strong out-of-distribution (OOD) generalisation by 18.7% higher success rate than VLA-Adapter under varying corruptions. However, the deeper blocks of the action expert, referred to as the expert head, remain unmergeable due to their strong task specialization. Consequently, each task keeps its own expert head, which in practice is typically the final block $L$. To address the challenge of mixed-task evaluation [knots, emr-merging], where the task identity is unknown at inference time, we adopt a training-free test-time task router that identifies the most relevant task directly from the input features. For each candidate task, the router runs the VLM with its corresponding task mask to obtain its hidden states, and then measures its response against principal components extracted from the value projections of a shared merged action expert. It selects the task with the highest response and activates the associated task mask and expert head. This allows MergeVLA to generalize to different tasks without additional supervision, enabling a single merged model to adaptively activate the right skill components for execution.
@@
Extensive experiments show that MergeVLA achieves success rates of $\mathbf{90.2\%}$, $\mathbf{72.2\%}$, and $\mathbf{70.7\%}$ on the LIBERO [libero], LIBERO-Plus [liberoplus], and RoboTwin [robotwin] benchmarks under the mixed-task evaluation setting, and $\mathbf{90.0\%}$ on real-world experiments with the SO101 robotic arm, demonstrating its effectiveness and robustness across cross-skill, cross-environment, and cross-embodiment evaluations. These results show that model merging is not only feasible for VLAs, but can serve as a scalable path toward generalist embodied agents.
@@ 2.1 Vision-Language-Action Models
Recent VLA models leverage the rich commonsense knowledge of large-scale VLMs and are fine-tuned on large-scale robotic trajectory data [bridgedatav2, OpenXEmbodiment, rt1, rt2] in an end-to-end manner to obtain more generalizable visuomotor policies [rt1, rt2, bridgedatav2, bridgedata, mtopt, qtopt, roboagent]. Among them, OpenVLA [OpenVLA] is a widely used open-source model based on Prismatic-7B [prismatic], which discretizes robot actions into rarely used vocabulary tokens and generates them autoregressively. Similarly, $\pi_{0}$ [pi0] builds on existing VLMs with a dual-system VLA architecture, introducing a lightweight action expert that uses conditional flow matching to generate continuous action chunks for faster inference. Building on this, $\pi_{0.5}$ [pi0.5] retains the text-generation ability of the VLM, enabling simultaneous high-level subtask descriptions and low-level action prediction. Recently, VLA-Adapter [vla-adapter] proposed a 0.5B-parameter dual-system model that coordinates the VLM and action expert through a novel mechanism, employing chunk-wise autoregressive generation for faster inference and strong performance at a small scale. While existing VLAs perform well on widely used robotic benchmarks [calvin, libero, robotwin, maniskill], their multi-task ability relies on joint training [OpenVLA, pi0, pi0.5, vla-adapter], making them inefficient. Besides, their tightly coupled dual systems [hirt, openhelix, dual1, pi0, vla-adapter] also hinder model merging.
@@ 2.2 Model Merging
Model merging enables efficient knowledge reuse by combining existing model weights to construct a unified model capable of performing multiple tasks. Early studies [WA1, WA2, modelsoup, regmean, fishermerging] showed that simple weighted averaging of checkpoints can improve performance and introduce multi-task capability. Task Arithmetic (TA) [TA] treats the parameter difference between fine-tuned and pretrained models as a task vector and merges them to integrate task-specific knowledge. Nonetheless, it overlooks various conflicts that may arise among different task vectors. Subsequent methods [ties, dare, breadcrumbs, pcb-merging, cabs, cat-merging, wudi] handle conflicts via parameter pruning or rescaling, whereas subspace-based ones [tsvm, knots, iso-cts, twin-merging] apply low-rank decomposition (e.g., SVD) for more consistent merging. Other studies [pem-composition, robust-merging, knots, do-merging] design merging methods specifically for parameter-efficient fine-tuning (PEFT) modules. In contrast to these pre-merge approaches, methods like [twin-merging, emr-merging, calm-merging, tall-mask, smile] perform test-time merging to handle severe task interference, substantially improving performance. Yet, little work has explored model merging in VLA models. The recent ReVLA [revla] applies merging to gradually reverse the vision backbone to mitigate visual catastrophic forgetting and enhance domain generalization, rather than to enable multi-task VLA capabilities. To bridge this gap, we propose MergeVLA which enables lightweight multi-task robotic learning.
@@ 3 Preliminary
VLA Architectures. Figure 1 illustrates the architectures of different VLAs. Some VLA models [OpenVLA, spatialvla, openvla-oft, tinyvla, magma] directly build upon existing VLMs, where only the language component is non-mergeable as refered to our experiments in Table 1. Others introduce an additional action expert [pi0, pi0.5, vla-adapter], among which VLA-Adapter trains its action expert from scratch. However, this tightly coupled dual-system design makes the overall model difficult to merge. MergeVLA simplifies this structure by removing the non-mergeable self-attention layers, enabling all components except the expert head to be merged effectively.
@@ 4 Our Approach: MergeVLA
Finetuning VLA models on individual manipulation tasks, while effective, produces isolated experts that cannot be trivially merged. As discussed in Section 5.2, applying standard model merging approaches to VLA specialists trained on LIBERO [libero] results in a complete breakdown, with all merged variants achieving zero success rate. This failure is unexpected given the success of merging strategies in language-only and vision-language domains, and suggests that VLA merging presents a more severe challenge. To understand this phenomenon, we conduct a systematic analysis of the parameter space and architectural behavior of mainstream VLAs. Our findings, as depicted in Figure 3, show that VLA unmergeability arises from two complementary failure modes:
@@
    Architectural Incompatibility of Action Experts: More critically, resolving LoRA interference is necessary but not sufficient. Even when the VLM is perfectly merged, simply averaging the action experts in architectures such as VLA-Adapter [vla-adapter] still yields 0% success. The right plot explains why: although the shallow blocks remain moderately aligned across tasks, the parameter distance explodes in the final layers. Such divergence arises because the action expert is trained entirely from scratch and contains self-attention layers that accumulate task-specific differences over depth. Instead of providing modular transformations, these layers propagate and amplify task-dependent signals, causing deeper blocks to become pathologically specialized to individual tasks. As a result, their parameters are inherently irreconcilable under any merging scheme.
Additionally, prior merging studies [knots, emr-merging] also note that the mixed-task setting is considerably more challenging than per-task evaluation, where the task identity is unknown. In practice, many approaches rely on hand-picked priors, e.g., task identity or an expert-specific prompt, to perform well, and degrade rapidly without them. Motivated by these insights, we therefore structure MergeVLA as a unified framework that addresses these challenges by: (1) Stabilizing VLM merging via task-specific masking (Sec 4.1) to address extreme LoRA parameter conflict (Q1); (2) redesigning the action expert (Sec 4.2) to mitigate architectural incompatibility (Q2); (3) introducing a learning-free task routing mechanism (Sec 4.3) to enable operation without a task identity at test time (Q3).
@@ 4.1 Task Conflicts in LoRA-Finetuned VLM
To address Q1, we first consider a standard LoRA update. Let $\Theta_{0}$ denote the pretrained checkpoint and $\Theta_{m}$ the LoRA-finetuned weights after finetuning on task $m\in\{1,\dots,M\}$. The task vector for task $m$ is defined as $\tau_{m}=\Theta_{m}-\Theta_{0}$. Most data-free $\mathcal{R}(\cdot)$ merging [TA, ties, emr-merging, tsvm, knots] then construct a single merged update $\tau_{\mathrm{merge}}$:
@@
where $\mathbb{I}[\cdot]$ is the indicator function that returns $1$ if the condition is true and 0 otherwise, and mask ratio $\lambda$ controls the tolerance for disagreement. This principled approach of pruning parameters based on their alignment consistency, adapted here for LoRA merging, shares its core formulation with task-vector compression methods [tall-mask].
@@
Empirically, as Figure 3 (left) shows, this selfish ratio on LIBERO for two merging methods as the number of merged tasks increases from 2 to 4. In all cases, $\text{ratio}_{\mathrm{selfish}}$ rises at around 75%, indicating that most parameters are selfish (kept exclusively by a single task) and underscoring the importance of task-specific masking to mitigate cross-task interference. Notably, applying the task mask encourages some LoRA-finetuned parameters to revert toward their pretrained weights, improving merge stability. A similar effect was observed in ReVLA [revla], where VLA finetuning was shown to overwrite pretrained visual knowledge. This forgetting becomes more pronounced when multiple experts are merged, as conflicting task updates distort the shared feature space. By sparsely activating merged LoRA parameters through the proposed mask, MergeVLA preserves pretrained visual–language representations and mitigates cross-task conflicts, enabling more consistent and stable merging across tasks.
Figure 3: Left: Selfish ratio of the masks from TA [TA] and TIES [ties] by merging different numbers of tasks. The selfish ratio is computed following Equation 4. Right: The average relative L2 distance across blocks between all pairs of action experts.
@@ 4.2 Redesigning the Action Expert for Mergeability
Our diagnosis proved that stabilizing the VLM (Q1) is insufficient; the action expert itself is a fundamental barrier (Q2). As shown in Section 5.2, applying only the LoRA task mask still results in 0% success rate. Our analysis pinpoints the incompatibility with the VLA-Adapter [vla-adapter] architecture (Fig. 1), which consists of $L$ transformer blocks trained from scratch. Each block contains a self-attention layer on the block input $\mathbf{x}^{i}$, two cross-attention layers conditioned on the VLM-provided hidden states $\mathbf{h}_{\text{T}}^{i}$ (task) and $\mathbf{h}_{\text{A}}^{i}$ (action), and a feed-forward network (FFN). A gating function is applied to the task stream, i.e., $\hat{\mathbf{h}}_{\text{T}}^{i}=g(\mathbf{h}_{\text{T}}^{i})$, with $g(\cdot)$ implemented as $\tanh$.
@@
These two architectural changes alone are remarkably effective, achieving an 18.7% higher success rate on the out-of-distribution (OOD) LIBERO-Plus [liberoplus] testbed. This confirms our new design better leverages the VLM’s robustness and is inherently more generalizable.
@@ 4.3 Test-time Task Routing
When the task identity is known, MergeVLA can already accomplish each skill by manually selecting the corresponding task mask $\mathbf{S}_{m}$ and expert head $\mathbf{H}_{m}^{l\rightarrow L}$. However, to operate without a known task identity (Q3) at inference time, the model must dynamically select these components based solely on the input observations, enabling cross-skill ability at a joint-task level [knots]. To this end, we propose a test-time task routing mechanism that infers task relevance directly from the model’s internal parameter subspaces, inspired by recent observations that fine-tuning pushes different tasks into distinguishable parameter subspaces [smile]. Given the merged LoRA parameters $\tau_{\mathrm{merge}}$ and task masks $\{\mathbf{S}_{m}\}_{m=1}^{M}$, we obtain $M$ masked VLM variants by applying each mask to the merged weights:
@@
Each masked VLM produces hidden states $\left[\mathbf{h}_{\mathrm{T}}^{l-1},\mathbf{h}_{\mathrm{A}}^{l-1}\right]$ from block (l1)(l{-}1), which is then forwarded to the corresponding (l1)(l{-}1)-th block of the merged action expert. Inside this block, two cross-attention paths are present: one conditioned on the task hidden state $\mathbf{h}_{\mathrm{T}}^{\,l-1}$ with parameters $(\mathbf{Q}_{\mathrm{T}}^{\,l-1},\mathbf{K}_{\mathrm{T}}^{\,l-1},\mathbf{V}_{\mathrm{T}}^{\,l-1})$, and the other conditioned on the action hidden state $\mathbf{h}_{\mathrm{A}}^{\,l-1}$ with $(\mathbf{Q}_{\mathrm{A}}^{\,l-1},\mathbf{K}_{\mathrm{A}}^{\,l-1},\mathbf{V}_{\mathrm{A}}^{\,l-1})$.
@@
A key design choice is which subspace to use for routing. We hypothesize that query $\mathbf{Q}$ and $\mathbf{K}$ govern attentional selection, thus being sensitive to input scaling and risk collapsing into task-specific subspaces. Empirically, value-based subspaces provide more stable and discriminative signals for routing, as they directly encode the task-dependent information written into the hidden states. We therefore analyze the principal components of the value projection matrices V of the (l1)(l{-}1)-th block for both paths via singular value decomposition (SVD):
$\mathbf{V}_{\mathrm{T}}^{\,l-1}$ $=\mathbf{L}_{\mathrm{T}}^{\,l-1}\,\mathbf{\Sigma}_{\mathrm{T}}^{\,l-1}\,(\mathbf{R}_{\mathrm{T}}^{\,l-1})^{\top},$ (6)
$\mathbf{V}_{\mathrm{A}}^{\,l-1}$ $=\mathbf{L}_{\mathrm{A}}^{\,l-1}\,\mathbf{\Sigma}_{\mathrm{A}}^{\,l-1}\,(\mathbf{R}_{\mathrm{A}}^{\,l-1})^{\top}.$
@@
We retain the top-$k_{r}$ right singular vectors in each path to form the dominant content components of the expert subspace: $\mathbf{P}_{\mathrm{T}}^{\,l-1}\in\mathbb{R}^{k\times d},$ $\mathbf{P}_{\mathrm{A}}^{\,l-1}\in\mathbb{R}^{k\times d},$ formed by taking the first $k_{r}$ rows of $(\mathbf{R}_{\mathrm{T}}^{\,l-1})^{\top}$ and $(\mathbf{R}_{\mathrm{A}}^{\,l-1})^{\top}$, respectively. For each task $m$, the hidden state from its masked VLM is projected onto these two subspaces to measure its activation strength,
$r_{\mathrm{T},m}=\big\|\mathbf{P}_{\mathrm{T}}^{\,l-1}\mathbf{h}^{\,l-1}_{\mathrm{A},m}\big\|_{2},r_{\mathrm{A},m}=\big\|\mathbf{P}_{\mathrm{A}}^{\,l-1}\mathbf{h}^{\,l-1}_{\mathrm{T},m}\big\|_{2}.$ (7)
Let $\mathbf{r}_{\mathrm{T}},\mathbf{r}_{\mathrm{A}}\in\mathbb{R}^{M}$ collect the scores $\{r_{\mathrm{T},m}\}_{m=1}^{M}$ and $\{r_{\mathrm{A},m}\}_{m=1}^{M}$, respectively. Here, the combined score vector is given by $\mathbf{r}=\tfrac{1}{2}\big(\mathbf{r}_{\mathrm{T}}+\mathbf{r}_{\mathrm{A}}\big)$.
The routing probabilities are calculated by softmax $p_{m}=\exp(\mathbf{r}_{m})/\sum_{j=1}^{M}\exp(\mathbf{r}_{j}).$
@@
Then we can use $\arg\max_{m}\,p_{m}$ to choose the task index $m^{*}$. Once $m^{*}$ is determined, the model uses the corresponding task mask $\mathbf{S}_{m^{*}}$ and expert head $\text{H}^{l\rightarrow L}_{m^{*}}$ to perform the forward pass. This design allows the router to infer the underlying task purely from the input-driven hidden representations of the masked VLMs, without any additional training or supervision. In practice, we find that a single routing step using the initial observation at t=0t{=}0 is sufficient to identify the correct task. The selected task mask and expert head are then fixed for the rest of the episode, avoiding repeated routing during inference. Although this mechanism requires maintaining $M$ task masks and their corresponding action heads, the additional computational and parameter overhead is minimal.

### v2

@@ 1 Introduction
Vision-Language-Action (VLA) models [28, 3, 27, 22, 4, 58, 2, 36, 38, 18, 46, 51, 19] have recently enabled robot agents to perform complex manipulation tasks by fine-tuning large vision-language models (VLMs) with millions of robotic demonstrations. By reformulating action learning as a language generation or policy decoding problem, VLA models inherit broad visual grounding and semantic understanding from VLMs, and have shown notable performance in single-task or single-embodiment settings. However, real-world generalist agents must support multiple skills, embodiments, and environments, requiring the ability to consolidate many independently fine-tuned VLA models into a single unified policy.
@@
A natural approach is model merging, which has proven effective in large language and vision models [21, 48, 31, 8, 53, 20, 39, 54, 56]. Those techniques can integrate multiple specialized models without joint retraining or revisiting their original datasets. Yet, when these approaches are applied to VLA experts fine-tuned on distinct manipulation tasks, the merged model exhibits near-zero success rate. This failure suggests that VLA finetuning induces structural specialization incompatible across tasks, which is rarely observed in conventional VLM merging.
@@
To uncover the root causes, as shown in Figure 1, we perform a fine-grained decomposition of trainable parameters across representative VLA architectures and identify two major sources of non-mergeability: First, LoRA updates within the VLM backbone diverge sharply across tasks. Each manipulation task reshapes the pretrained representation space to satisfy its own perception-control alignment. Direct averaging  [21, 39] or sign-resolved merging [48] thus reactivates irrelevant or even contradictory parameters, corrupting shared vision-language subspaces and degrading task-invariant semantics. Second, the train-from-scratch action decoders accumulate strong task-specific dependencies across blocks through self-attention feedback. This coupling spreads localized task information globally, breaking modularity and preventing compositional merging even under identical architectures and initialization.
@@
Built upon these findings, we introduce MergeVLA, a merge–oriented VLA architecture that preserves mergeability by design. When executing a specific task, MergeVLA applies sparsely activated LoRA adapters, implemented via task masks, to selectively activate the merged parameters contributing to task-relevant responses while suppressing those that mislead other tasks. Moreover, MergeVLA reconfigures the action expert to remove self-attention propagation, relying solely on cross-attention pathways. This eliminates the sustained accumulation of task dependence across blocks, allowing most layers to be effectively merged using simple weight averaging. Due to these architectural modifications, it surprisingly shows strong out-of-distribution (OOD) generalisation by 13.4% higher success rate than VLA-Adapter under varying corruptions. However, the deeper blocks of the action expert, referred to as the expert head, remain unmergeable due to their strong task specialization. Consequently, each task keeps its own expert head, which in practice is typically the final block $L$. To address the challenge of mixed-task evaluation [39, 20], where the task identity is unknown at inference time, we adopt a training-free test-time task router that identifies the most relevant task directly from the input features. For each candidate task, the router runs the VLM with its corresponding task mask to obtain its hidden states, and then measures its response against principal components extracted from the value projections of a shared merged action expert. It selects the task with the highest response and activates the associated task mask and expert head. This allows MergeVLA to generalize to different tasks without additional supervision, enabling a single merged model to adaptively activate the right skill components for execution.
@@
Extensive experiments show that MergeVLA achieves success rates of $\mathbf{90.2\%}$, $\mathbf{62.5\%}$, and $\mathbf{70.7\%}$ on the LIBERO [29], LIBERO-Plus [15], and RoboTwin [7] benchmarks under the mixed-task evaluation setting, and $\mathbf{90.0\%}$ on real-world experiments with the SO101 robotic arm, demonstrating its effectiveness and robustness across cross-skill, cross-environment, and cross-embodiment evaluations. These results show that model merging is not only feasible for VLAs, but can serve as a scalable path toward generalist embodied agents.
@@ 2.1 Vision-Language-Action Models
Recent VLA models leverage the rich commonsense knowledge of large-scale VLMs and are fine-tuned on large-scale robotic trajectory data [43, 35, 4, 58] in an end-to-end manner to obtain more generalizable visuomotor policies [4, 58, 43, 14, 25, 24, 1]. Among them, OpenVLA [28] is a widely used open-source model based on Prismatic-7B [26], which discretizes robot actions into rarely used vocabulary tokens and generates them autoregressively. Similarly, $\pi_{0}$ [3] builds on existing VLMs with a dual-system VLA architecture, introducing a lightweight action expert that uses conditional flow matching to generate continuous action chunks for faster inference. Building on this, $\pi_{0.5}$ [22] retains the text-generation ability of the VLM, enabling simultaneous high-level subtask descriptions and low-level action prediction. Recently, VLA-Adapter [45] proposed a 0.5B-parameter dual-system model that coordinates the VLM and action expert through a novel mechanism, employing chunk-wise autoregressive generation for faster inference and strong performance at a small scale. While existing VLAs perform well on widely used robotic benchmarks [33, 29, 7, 42], their multi-task ability relies on joint training [28, 3, 22, 45], making them inefficient. Besides, their tightly coupled dual systems [55, 9, 5, 3, 45] also hinder model merging.
@@ 2.2 Model Merging
Model merging enables efficient knowledge reuse by combining existing model weights to construct a unified model capable of performing multiple tasks. Early studies [16, 34, 47, 23, 32] showed that simple weighted averaging of checkpoints can improve performance and introduce multi-task capability. Task Arithmetic (TA) [21] treats the parameter difference between fine-tuned and pretrained models as a task vector and merges them to integrate task-specific knowledge. Nonetheless, it overlooks various conflicts that may arise among different task vectors. Subsequent methods [48, 53, 10, 12, 52, 40, 8] handle conflicts via parameter pruning or rescaling, whereas subspace-based ones [17, 39, 31, 30] apply low-rank decomposition (e.g., SVD) for more consistent merging. Other studies [56, 54, 39, 57] design merging methods specifically for parameter-efficient fine-tuning (PEFT) modules. In contrast to these pre-merge approaches, methods like [30, 20, 49, 44, 41] perform test-time merging to handle severe task interference, substantially improving performance. Yet, little work has explored model merging in VLA models. The recent ReVLA [11] applies merging to gradually reverse the vision backbone to mitigate visual catastrophic forgetting and enhance domain generalization, rather than to enable multi-task VLA capabilities. To bridge this gap, we propose MergeVLA which enables lightweight multi-task robotic learning.
@@ 3 Preliminary
VLA Architectures. Figure 1 illustrates the architectures of different VLAs. Some VLA models [28, 36, 27, 46, 51] directly build upon existing VLMs, where only the language component is non-mergeable as refered to our experiments in Table 1. Others introduce an additional action expert [3, 22, 45], among which VLA-Adapter trains its action expert from scratch. However, this tightly coupled dual-system design makes the overall model difficult to merge. MergeVLA simplifies this structure by removing the non-mergeable self-attention layers, enabling all components except the expert head to be merged effectively.
@@ 4 Our Approach: MergeVLA
Finetuning VLA models on individual manipulation tasks, while effective, produces isolated experts that cannot be trivially merged. As discussed in Section 5.2, applying standard model merging approaches to VLA specialists trained on LIBERO [29] results in a complete breakdown, with all merged variants achieving zero success rate. This failure is unexpected given the success of merging strategies in language-only and vision-language domains, and suggests that VLA merging presents a more severe challenge. To understand this phenomenon, we conduct a systematic analysis of the parameter space and architectural behavior of mainstream VLAs. Our findings, as depicted in Figure 3, show that VLA unmergeability arises from two complementary failure modes:
@@
    Architectural Incompatibility of Action Experts: More critically, resolving LoRA interference is necessary but not sufficient. Even when the VLM is perfectly merged, simply averaging the action experts in architectures such as VLA-Adapter [45] still yields 0% success. The right plot explains why: although the shallow blocks remain moderately aligned across tasks, the parameter distance explodes in the final layers. Such divergence arises because the action expert is trained entirely from scratch and contains self-attention layers that accumulate task-specific differences over depth. Instead of providing modular transformations, these layers propagate and amplify task-dependent signals, causing deeper blocks to become pathologically specialized to individual tasks. As a result, their parameters are inherently irreconcilable under any merging scheme.
Additionally, prior merging studies [39, 20] also note that the mixed-task setting is considerably more challenging than per-task evaluation, where the task identity is unknown. In practice, many approaches rely on hand-picked priors, e.g., task identity or an expert-specific prompt, to perform well, and degrade rapidly without them. Motivated by these insights, we therefore structure MergeVLA as a unified framework that addresses these challenges by: (1) Stabilizing VLM merging via task-specific masking (Sec 4.1) to address extreme LoRA parameter conflict (Q1); (2) redesigning the action expert (Sec 4.2) to mitigate architectural incompatibility (Q2); (3) introducing a learning-free task routing mechanism (Sec 4.3) to enable operation without a task identity at test time (Q3).
@@ 4.1 Task Conflicts in LoRA-Finetuned VLM
To address Q1, we first consider a standard LoRA update. Let $\Theta_{0}$ denote the pretrained checkpoint and $\Theta_{m}$ the LoRA-finetuned weights after finetuning on task $m\in\{1,\dots,M\}$. The task vector for task $m$ is defined as $\tau_{m}=\Theta_{m}-\Theta_{0}$. Most data-free methods $\mathcal{R}(\cdot)$ construct a single merged update $\tau_{\mathrm{merge}}$ by merging the updates [21, 48, 20, 17, 39]:
@@
where $\mathbb{I}[\cdot]$ is the indicator function that returns $1$ if the condition is true and 0 otherwise, and mask ratio $\lambda$ controls the tolerance for disagreement. This principled approach of pruning parameters based on their alignment consistency, adapted here for LoRA merging, shares its core formulation with task-vector compression methods [44].
@@
Empirically, as Figure 3 (left) shows, this selfish ratio on LIBERO for two merging methods as the number of merged tasks increases from 2 to 4. In all cases, $\text{ratio}_{\mathrm{selfish}}$ rises at around 75%, indicating that most parameters are selfish (kept exclusively by a single task) and underscoring the importance of task-specific masking to mitigate cross-task interference. Notably, applying the task mask encourages some LoRA-finetuned parameters to revert toward their pretrained weights, improving merge stability. A similar effect was observed in ReVLA [11], where VLA finetuning was shown to overwrite pretrained visual knowledge. This forgetting becomes more pronounced when multiple experts are merged, as conflicting task updates distort the shared feature space. By sparsely activating merged LoRA parameters through the proposed mask, MergeVLA preserves pretrained visual–language representations and mitigates cross-task conflicts, enabling more consistent and stable merging across tasks.
Figure 3: Left: Selfish ratio of the masks from TA [21] and TIES [48] by merging different numbers of tasks. The selfish ratio is computed following Equation 4. Right: The average relative L2 distance across blocks between all pairs of action experts.
@@ 4.2 Redesigning the Action Expert for Mergeability
Our diagnosis proved that stabilizing the VLM (Q1) is insufficient; the action expert itself is a fundamental barrier (Q2). As shown in Section 5.2, applying only the LoRA task mask still results in 0% success rate. Our analysis pinpoints the incompatibility with the VLA-Adapter [45] architecture (Fig. 1), which consists of $L$ transformer blocks trained from scratch. Each block contains a self-attention layer on the block input $\mathbf{x}^{i}$, two cross-attention layers conditioned on the VLM-provided hidden states $\mathbf{h}_{\text{T}}^{i}$ (task) and $\mathbf{h}_{\text{A}}^{i}$ (action), and a feed-forward network (FFN). A gating function is applied to the task stream, i.e., $\hat{\mathbf{h}}_{\text{T}}^{i}=g(\mathbf{h}_{\text{T}}^{i})$, with $g(\cdot)$ implemented as $\tanh$.
@@
These two architectural changes alone are remarkably effective, achieving an 13.4% higher success rate on the out-of-distribution (OOD) LIBERO-Plus [15] testbed. This confirms our new design better leverages the VLM’s robustness and is inherently more generalizable.
@@ 4.3 Test-time Task Routing
When the task identity is known, MergeVLA can already accomplish each skill by manually selecting the corresponding task mask $\mathbf{S}_{m}$ and expert head $\mathbf{H}_{m}^{l\rightarrow L}$. However, to operate without a known task identity (Q3) at inference time, the model must dynamically select these components based solely on the input observations, enabling cross-skill ability at a joint-task level [39]. To this end, we propose a test-time task routing mechanism that infers task relevance directly from the model’s internal parameter subspaces, inspired by recent observations that fine-tuning pushes different tasks into distinguishable parameter subspaces [41]. Given the merged LoRA parameters $\tau_{\mathrm{merge}}$ and task masks $\{\mathbf{S}_{m}\}_{m=1}^{M}$, we obtain $M$ masked VLM variants by applying each mask to the merged weights:
@@
Each masked VLM produces hidden states $\left[\mathbf{h}_{\mathrm{T}}^{l},\mathbf{h}_{\mathrm{A}}^{l}\right]$ from block $l$, which is then forwarded to the corresponding $l$-th block of the action expert. Inside this block, two cross-attention paths are present: one conditioned on the task hidden state $\mathbf{h}_{\mathrm{T}}^{\,l}$ with parameters $(\mathbf{Q}_{\mathrm{T}}^{\,l},\mathbf{K}_{\mathrm{T}}^{\,l},\mathbf{V}_{\mathrm{T}}^{\,l})$, and the other conditioned on the action hidden state $\mathbf{h}_{\mathrm{A}}^{\,l}$ with $(\mathbf{Q}_{\mathrm{A}}^{\,l},\mathbf{K}_{\mathrm{A}}^{\,l},\mathbf{V}_{\mathrm{A}}^{\,l})$.
@@
A key design choice is which subspace to use for routing. We hypothesize that query $\mathbf{Q}$ and $\mathbf{K}$ govern attentional selection, thus being sensitive to input scaling and risk collapsing into task-specific subspaces. Empirically, value-based subspaces provide more stable and discriminative signals for routing, as they directly encode the task-dependent information written into the hidden states. We therefore analyze the principal components of the value projection matrices V of the $l$-th block for both paths via singular value decomposition (SVD):
$\mathbf{V}_{\mathrm{T}}^{\,l}$ $=\mathbf{L}_{\mathrm{T}}^{\,l}\,\mathbf{\Sigma}_{\mathrm{T}}^{\,l}\,(\mathbf{R}_{\mathrm{T}}^{\,l})^{\top},$ (6)
$\mathbf{V}_{\mathrm{A}}^{\,l}$ $=\mathbf{L}_{\mathrm{A}}^{\,l}\,\mathbf{\Sigma}_{\mathrm{A}}^{\,l}\,(\mathbf{R}_{\mathrm{A}}^{\,l})^{\top}.$
@@
We retain the top-$k_{r}$ right singular vectors in each path to form the dominant content components of the expert subspace: $\mathbf{P}_{\mathrm{T}}^{\,l}\in\mathbb{R}^{k_{r}\times d},$ $\mathbf{P}_{\mathrm{A}}^{\,l}\in\mathbb{R}^{k_{r}\times d},$ formed by taking the first $k_{r}$ rows of $(\mathbf{R}_{\mathrm{T}}^{\,l})^{\top}$ and $(\mathbf{R}_{\mathrm{A}}^{\,l})^{\top}$, respectively. For each task $m$, the hidden state from its masked VLM is projected onto these two subspaces to measure its activation strength,
$r_{\mathrm{T},m}=\big\|\mathbf{P}_{\mathrm{T}}^{\,l}\mathbf{h}^{\,l}_{\mathrm{A},m}\big\|_{2},r_{\mathrm{A},m}=\big\|\mathbf{P}_{\mathrm{A}}^{\,l}\mathbf{h}^{\,l}_{\mathrm{T},m}\big\|_{2}.$ (7)
@@
Let $\mathbf{r}_{\mathrm{T}},\mathbf{r}_{\mathrm{A}}\in\mathbb{R}^{M}$ collect the scores $\{r_{\mathrm{T},m}\}_{m=1}^{M}$ and $\{r_{\mathrm{A},m}\}_{m=1}^{M}$, respectively. Here, the combined score vector is given by $\mathbf{r}=\tfrac{1}{2}\big(\mathbf{r}_{\mathrm{T}}+\mathbf{r}_{\mathrm{A}}\big)$. The routing probabilities are calculated by softmax $p_{m}=\operatorname{softmax}(\mathbf{r})_{m}.$ Then we can use $\arg\max_{m}\,p_{m}$ to choose the task index $m^{*}$. Once $m^{*}$ is determined, the model uses the corresponding task mask $\mathbf{S}_{m^{*}}$ and expert head $\text{H}^{l\rightarrow L}_{m^{*}}$ to perform the forward pass. This design allows the router to infer the underlying task purely from the input-driven hidden representations of the masked VLMs, without any additional training or supervision. In practice, we find that a single routing step using the initial observation at t=0t{=}0 is sufficient to identify the correct task. The selected task mask and expert head are then fixed for the rest of the episode, avoiding repeated routing during inference. Although this mechanism requires maintaining $M$ task masks and their corresponding action heads, the additional computational and parameter overhead is minimal.